Documenta Mathematica

Band 8, 2003

Elmar Grosse-Klönne

On Families of Pure Slope L-Functions  1–42

Bas Edixhoven, Chandrashekhar Khare

Hasse Invariant and Group Cohomology  43–50

Ivan Panin, Kirill Zainoulline

Variations on the Bloch-Ogus Theorem  51–67

Patrick Brosnan A Short Proof of Rost Nilpotence

via Refined Correspondences  69–78

Wolfgang Krieger and Kengo Matsumoto A Lambda-Graph System

for the Dyck Shift and Its K-Groups  79–96

G. van der Geer, T. Katsura On the Height of Calabi-Yau Varieties

in Positive Characteristic  97–113

Norbert Hoffmann Stability of Arakelov Bundles

and Tensor Products without Global Sections  115–123

Joost van Hamel The Reciprocity Obstruction for Rational Points on Compactifications of Torsors under Tori

over Fields with Global Duality  125–142

Michael Puschnigg Diffeotopy Functors of ind-Algebras

and Local Cyclic Cohomology  143–245

Jean-Paul Bonnet Un Isomorphisme Motivique entre Deux Variétés Homogènes Projectives

sous l’Action d’un Groupe de Type G_{2}  247–277

Grégory Berhuy, Giordano Favi Essential Dimension: a Functorial Point of View

(After A. Merkurjev)  279–330

Brian Conrad, Bas Edixhoven, William Stein

J_{1}(p) Has Connected Fibers  331–408

Bjørn Ian Dundas, Oliver Röndigs, Paul Arne Østvær

Enriched Functors and Stable Homotopy Theory  409–488

Bjørn Ian Dundas, Oliver Röndigs, Paul Arne Østvær

Motivic Functors  489–525

Daniel Krashen Severi-Brauer Varieties

of Semidirect Product Algebras  527–546

Rupert L. Frank On the scattering theory of the Laplacian with a periodic boundary condition.

I. Existence of wave operators  547–565

Douglas Bridges and Luminiţa Vîţă

Separatedness in Constructive Topology  567–576

Chin-Lung Wang

Curvature Properties of the Calabi-Yau Moduli  577–590

Marc A. Nieper-Wißkirchen Calculation of Rozansky-Witten Invariants on the Hilbert Schemes of Points on a K3 Surface

and the Generalised Kummer Varieties  591–623